a vertical wall 18 feet high casts a shadow that is 6 feet wide. If Joe is 5'3' tall, how far away from the wall can he stand and still be entirely in the shade?
Answer by kapilsinghi(68)

18f cast 6f shadow
5.25 feet will cast 5.25*6/18 = 5.25/3 = 1.75f
if second shadow is to be completely merged in the first one then the person should be standing 6 - 1.75 feet from wall
= 4.25 feet from wall
kapilsinghi123@gmail.com